# Randomization Tests for Conditional Group Symmetry

### v1

@@ B.4 Proofs for Section 6
where $F_{B,p(\hat{\Delta}_{n})}$ is the $\textrm{Binomial}(B,p(\hat{\Delta}_{n}))$ distribution function. It is known that for $q\geq p$, a $\textrm{Binomial}(B,q)$-distributed random variable $X_{B,q}$ stochastically dominates a $\textrm{Binomial}(B,p)$-distributed random variable $X_{B,p}$, in the sense that
$\mathbb{P}(X_{B,q}\geq k)\geq\mathbb{P}(X_{B,p}\geq k)\;,\quad\text{all }k\in[B]\;.$
This implies the reverse ordering for the CDFs, so that the bound $p(\hat{\Delta}_{n})\leq\hat{p}(\hat{\Delta}_{n})$ implies
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n},\hat{\Delta}_{n})\geq F_{B,\hat{p}(\hat{\Delta}_{n})}(\lfloor\alpha(B+1)-1\rfloor)$
$=\sum_{k=0}^{\lfloor\alpha(B+1)-1\rfloor}\binom{B}{k}\hat{p}(\hat{\Delta}_{n})^{k}(1-\hat{p}(\hat{\Delta}_{n}))^{B-k}\;.$
Now, applying Lemma B.5 to (27) with $\hat{P}_{1:n},Q_{1:n}$ the empirical distributions of the observations $Z_{1:n}$ and the randomizations $Z_{1:n}^{(0)}$, respectively, conditioned on $S_{1:n}$, we have that with probability at least $1-e^{-\eta}$,
$\left|T(\hat{P}_{1:n},\hat{P}^{(0)}_{n})-T(P_{1:n},P^{(0)}_{1:n})\right|\leq 2\nu\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{2\frac{\nu^{2}}{n}+32\nu^{2}\sqrt{\frac{2}{n^{3}}}}+\frac{8\eta\nu}{3n}\;.$
Therefore, letting $\Delta_{n}\coloneq T(P_{1:n},P^{(0)}_{1:n})$ (as in (28)), with probability at least $1-e^{-\eta}$,
$\hat{\Delta}_{n}\geq\Delta_{n}-2\nu\sqrt{\frac{2}{n}}-\sqrt{\eta}\sqrt{2\frac{\nu^{2}}{n}+32\nu^{2}\sqrt{\frac{2}{n^{3}}}}-\frac{8\eta\nu}{3n}\;,$
Putting this all together, with probability at least $1-e^{-\eta}$,
$\mathbb{P}(p_{B}\leq\alpha$ $\mid S_{1:n})\geq\sum_{k=0}^{\lfloor\alpha(B+1)-1\rfloor}\binom{B}{k}p_{\eta}(\Delta_{n})^{k}(1-p_{\eta}(\Delta_{n}))^{B-k}\;,$
where
$p_{\eta}(\Delta_{n})=\exp\left(-\frac{3}{2}\Delta_{n}+R_{n}(\eta)\right)\;,$
with
$R_{n}(\eta)\coloneq\frac{9}{2n}\nu^{2}+\frac{16\nu^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu\sqrt{\frac{2}{n}}+\frac{3\nu}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu}{n}\;.$
B.4.2 Proof of Corollary 6.2
Proof of Corollary 6.2.
Let $B_{\eta}(\Delta_{n})=Bp_{\eta}(\Delta_{n})$. By the Chernoff bound for binomial random variables,
$\mathbb{P}(N_{B}\leq B_{\alpha}\mid S_{1:n})\geq 1-\exp\left(-B_{\alpha}\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}+B_{\alpha}-B_{\eta}(\Delta_{n})\right)\;.$
This is at least $\beta$ if and only if
$-B_{\alpha}\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}+B_{\alpha}-B_{\eta}(\Delta_{n})\leq\ln(1-\beta)\;.$ (30)
Now,
$\ln\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}=-\ln\left(1-\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)=\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}\;,$
and therefore (30) becomes
$B_{\alpha}\sum_{k=1}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}-(B_{\alpha}-B_{\eta}(\Delta_{n}))$
$=B_{\alpha}\sum_{k=2}^{\infty}\frac{1}{k}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{k}\geq-\ln(1-\beta)\;.$
The inequality is satisfied if
$B_{\alpha}\frac{1}{2}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{2}\geq-\ln(1-\beta)\;.$
Solving for $\Delta_{n}$ yields
$\Delta_{n}\geq\frac{2}{3}\left(R_{n}(\eta)-\ln\frac{B_{\alpha}}{B}-\ln\left(1-\sqrt{\frac{-2\ln(1-\beta)}{B_{\alpha}}}\right)\right)\;.$
B.5 Power lower bounds for adaptive statistics
@@
$\bar{\Delta}_{n}\coloneq\frac{1}{L}\sum_{\ell=1}^{L}\sqrt{\mathbb{E}\left[\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z^{\prime(0)}_{1:n})\mid S_{1:n}\right]}\;.$
For convenience, let $R_{n}^{\vee}(\eta)$ denote the maximum over $\ell\in[L]$ of $R_{n,\ell}(\eta)$, which is obtained by substituting $\nu_{\ell}$ for $\nu$ into (19).
Proposition B.7.
Let $T$ be the FUSE statistic using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. Assume that for each $\ell\in\{1,\dotsc,L\}$, $\Delta_{n,\ell}\geq 2\nu_{\ell}\sqrt{2/n}$. Then for every $\eta>0$, for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $1-Le^{-\eta}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
@@
$p_{\eta}(\bar{\Delta}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+R_{n}^{\vee}(\eta)\right)\right\}\;.$
@@
Using this, a lower bound for $\bar{\Delta}_{n}$ that guarantees power $\beta$ with high probability, analogous to the one in Corollary 6.2, can be obtained.
For the following proof, define $\nu_{\vee}\coloneq\sup_{\ell}\nu_{\ell}$, so that
$a^{\vee}\coloneq\sup_{\ell\in[L]}a_{\ell}\leq\frac{\nu_{\vee}^{2}}{n}\left(1+16\sqrt{\frac{2}{n}}\right)\;,$
and similarly $\bar{H}^{\vee}\coloneq\sup_{\ell}\bar{H}_{\ell}\leq 2\nu_{\vee}\sqrt{2/n}$. Finally, let
$R_{n}^{\vee}(\eta)\coloneq\frac{9}{2n}\nu_{\vee}^{2}+\frac{16\nu_{\vee}^{2}\sqrt{2}}{\sqrt{n^{3}}}+6\nu_{\vee}\sqrt{\frac{2}{n}}+\frac{3\nu_{\vee}}{2}\sqrt{\eta}\sqrt{\frac{2}{n}+32\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta\nu_{\vee}}{n}\;.$
Proof of Proposition B.7.
Using the V-statistic to estimate MMD, recall that
@@
$\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)\;.$
Therefore, using the fact that if $\mathbb{E}[X_{1}]\geq\mathbb{E}[X_{2}]$ then $\sup X_{1}\geq\mathbb{E}[X_{2}]$, denoting $\mathcal{A}_{1:n}=(S_{1:n},Z_{1:n},Z^{(0)}_{1:n})$,
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\mathbb{P}\left(\sup_{\ell}\exp\left[\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\right]\geq\right.$
$\qquad\qquad\left.\frac{1}{L}\sum_{\ell=1}^{L}\exp\left[\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},{Z}_{1:n}^{(0)})\right]\mid\mathcal{A}_{1:n}\right)$
@@
$=\mathbb{P}\left(\sup_{\ell}\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)\;.$
By a Fréchet inequality and Lemma B.5,
@@
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\bigg{|}\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}\mid\mathcal{A}_{1:n}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})}{3a_{\ell}+\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})^{1/2}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Now by Jensen’s inequality,
$\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)$
$\geq\frac{1}{\omega}\frac{1}{L}\sum_{\ell=1}^{L}\log\exp\left(\omega\widehat{\text{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)$
$\geq\left(\frac{1}{L}\sum_{\ell=1}^{L}\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\bigg{|}\right)^{2}\;.$
Therefore,
@@
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\frac{1}{L}\sum_{\ell=1}^{L}\bigg{|}\widehat{\text{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},{Z}_{1:n}^{(0)})\bigg{|}+\frac{9}{2}a^{\vee}+\frac{3}{2}\bar{H}^{\vee}\right)\right\}\;.$
Thus, for any $\eta>0$, with probability at least $1-Le^{-\eta}$,
@@
$\mathbb{P}\left(\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n}^{(b)},{Z^{\prime}}_{1:n}^{(b)})\geq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},{Z}_{1:n}^{(0)})\mid\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+R^{\vee}_{n}(\eta)\right)\right\}\;.$
The rest of the proof is analogous to the proof of Theorem 6.1. ∎
@@
$\Delta_{n}^{\vee}=\sup_{\ell\in[L]}\Delta_{n,\ell}\coloneq\sup_{\ell\in[L]}\sqrt{\mathbb{E}\left[\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:{n}},{Z}_{1:{n}}^{(0)})\mid S_{1:n}\right]}$

### v2

@@ C.4 Additional results and proofs for Section 5
[…]
$\mathbb{P}(X_{B,q}\geq\ell)\geq\mathbb{P}(X_{B,p}\geq\ell)\;.$
The inequality is reversed for the CDFs, so that the bound $p(\hat{\Delta}_{n})\leq p_{\eta}(\Delta_{n})$ implies
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})$ $\geq F_{B,p_{\eta}(\Delta_{n})}(\lfloor\alpha(B+1)-1\rfloor)\;.$
Thus, we have showed that with probability at least $1-e^{-\eta}$,
$\mathbb{P}(p_{B}\leq\alpha\mid S_{1:n})\geq\sum_{\ell=0}^{\lfloor\alpha(B+1)-1\rfloor}\binom{B}{\ell}p_{\eta}(\Delta_{n})^{\ell}(1-p_{\eta}(\Delta_{n}))^{B-\ell}\;.$
C.4.2 Lower bound for high-probability rejection of alternatives
Using the Chernoff bound on the upper tail of deviations of a binomial random variable, we can obtain an explicit lower bound on values of $\Delta_{n}$ that correspond to alternatives for which the test (with high probability) rejects the null with probability at least $1-\beta$ for $\beta\in(0,1)$. For convenience, let $B_{\alpha}\coloneq\lfloor\alpha(B+1)-1\rfloor$.
Corollary C.8.
$\Delta_{n}\geq\frac{2}{3}\left(C_{n}(\eta)-\ln\frac{B_{\alpha}}{B}-\ln\left(1-\sqrt{\frac{-2\ln\beta}{B_{\alpha}}}\right)\right)\;,$
then for almost every sequence $(S_{i})_{i\geq 1}$, conditionally on $S_{1:n}$, we have that $\mathbb{P}(p_{B}\leq\alpha|S_{1:n})\geq 1-\beta$ with probability over $(Z_{1:n},{Z}^{(0)}_{1:n})$ at least $1-e^{-\eta}$.
In the limit of $B,n\to\infty$, we see that if $\text{MMD}(P,P^{(0)})\geq-\frac{2}{3}\ln\alpha$, then the test will reject with probability at least $1-\beta$.
Proof of Corollary C.8.
Let $B_{\eta}(\Delta_{n})\coloneqq Bp_{\eta}(\Delta_{n})$. Applying a Chernoff bound on the upper tail of deviations of a binomial random variable to $N_{B}\sim\mathrm{Binomial}(B,p_{\eta}(\Delta_{n}))$,
$\mathbb{P}(N_{B}\geq B_{\alpha}\mid S_{1:n})$ $=\mathbb{P}\left(\left.N_{B}\geq\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\right)B_{\eta}(\Delta_{n})\;\right|\;S_{1:n}\right)$
$\leq\exp\left\{-B_{\eta}(\Delta_{n})\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\ln\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\right)-\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}+1\right)\right\}$
$=\exp\left(-B_{\alpha}\ln\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\right)+B_{\alpha}-B_{\eta}(\Delta_{n})\right)\;.$
This bound is at most $\beta$ if and only if
$B_{\alpha}\ln\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\right)-(B_{\alpha}-B_{\eta}(\Delta_{n}))\geq-\ln\beta\;.$
Note that
$\ln\left(\frac{B_{\alpha}}{B_{\eta}(\Delta_{n})}\right)=-\ln\left(1-\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)=\sum_{\ell=1}^{\infty}\frac{1}{\ell}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{\ell}\;,$
where the last equality follows from the Taylor expansion of $\ln(1-x)$ for $|x|<1$, and so an equivalent condition is
$B_{\alpha}\sum_{\ell=2}^{\infty}\frac{1}{\ell}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{\ell}\geq-\ln\beta\;.$
The condition is satisfied if
$B_{\alpha}\frac{1}{2}\left(\frac{B_{\alpha}-B_{\eta}(\Delta_{n})}{B_{\alpha}}\right)^{2}\geq-\ln\beta\;,$
or equivalently,
$B_{\eta}(\Delta_{n})\leq B_{\alpha}\left(1-\sqrt{-\frac{2\ln\beta}{B_{\alpha}}}\right)\;.$
Substituting in the definition for $B_{\eta}(\Delta_{n})$ and rearranging for $\Delta_{n}$ yields
$\Delta_{n}\geq\frac{2}{3}\left(C_{n}(\eta)-\ln\left(\frac{B_{\alpha}}{B}\right)-\ln\left(1-\sqrt{-\frac{2\ln\beta}{B_{\alpha}}}\right)\right)\;.$
C.5 Power lower bounds for adaptive statistics
@@
$\bar{\Delta}_{n}\coloneqq\frac{1}{L}\sum_{\ell=1}^{L}\sqrt{\mathbb{E}\left[\left.\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z^{(0)}_{1:n})\;\right|\;S_{1:n}\right]}\;.$
For convenience, let $C_{n}^{\vee}(\eta)$ denote the maximum over $\ell\in[L]$ of $C_{n,\ell}(\eta)$, which is obtained by substituting $\kappa_{\ell}$ for $\kappa$ into (14). Define $\kappa_{\lor}\coloneqq\sup_{\ell}\kappa_{\ell}$.
Proposition C.9.
Let $T$ be the FUSE statistic using $\widehat{\textrm{MMD}}^{2}_{v,\sigma_{\ell}}$ and exact randomization in the CRT. For any $\eta>0$, if
$\bar{\Delta}_{n}>2\kappa_{\lor}\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\;,$ (24)
then for almost every sequence $S_{1:n}$, conditionally on $S_{1:n}$, with probability at least $\max\{0,1-Le^{-\eta}\}$ over $(Z_{1:n},{Z}^{(0)}_{1:n})$,
@@
$p_{\eta}(\bar{\Delta}_{n})=\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+C_{n}^{\vee}(\eta)\right)\right\}\;.$
@@
Using this, a lower bound for $\bar{\Delta}_{n}$ that guarantees power $1-\beta$ with high probability, analogous to the one in Corollary C.8, can be obtained.
For the following proof, define
$a^{\lor}\coloneqq\sup_{\ell\in[L]}a_{\ell}\leq\kappa_{\lor}^{2}\left(\frac{2}{n}+16\sqrt{\frac{2}{n^{3}}}\right)$ (25)
and similarly
$\xi_{n}^{\lor}\coloneqq\sup_{\ell\in[L]}\xi_{n,\ell}\leq 2\kappa_{\lor}\sqrt{\frac{2}{n}}\;.$ (26)
Proof of Proposition C.9.
Let
$T(Z_{1:n},Z_{1:n}^{\prime})$ $\coloneqq\widehat{\mathrm{FUSE}}_{\omega}(Z_{1:n},Z_{1:n}^{\prime})$
@@
$=\frac{1}{\omega}\log\left(\frac{1}{L}\sum_{\ell=1}^{L}\exp\left(\omega\widehat{\mathrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{\prime})\right)\right)\;,$
and let $\widehat{T}\coloneqq T(Z_{1:n},Z_{1:n}^{(0)})$ denote the observed test statistic. Denote $\mathcal{A}_{1:n}\coloneqq(S_{1:n},Z_{1:n},Z_{1:n}^{(0)})$. Observe that
[…]
@@
$\leq\mathbb{P}\left\{\left.\frac{1}{\omega}\log\left(\sup_{\ell\in[L]}\exp\left(\omega\widehat{\mathrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right)\right)\geq\widehat{T}\;\right|\;\mathcal{A}_{1:n}\right\}$
@@
$=\mathbb{P}\left(\left.\sup_{\ell\in[L]}\widehat{\mathrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\geq\widehat{T}\;\right|\;\mathcal{A}_{1:n}\right)$
@@
$\leq\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|\geq\sqrt{\widehat{T}}\;\right|\;\mathcal{A}_{1:n}\right)$
@@
$=\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n}\right)$
$\leq\min\left\{1,\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right$.
$\qquad\qquad+\mathbbm{1}\left\{\sqrt{\widehat{T}}\leq\xi_{n}^{\lor}\right\}\Bigg\}\;,$
where the last inequality follows from the same argument as that in Theorem 5.1. We first derive a high-probability upper bound for the first term. Applying a Fréchet inequality (first inequality), Lemma C.6 (third inequality), the same factoring trick as in (22) (fourth inequality), and Jensen’s inequality (last two inequalities), we obtain
$p^{*}$ $\coloneqq\mathbb{P}\left(\left.\sup_{\ell\in[L]}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left.\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n}^{\lor}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\mathbb{P}\left(\left.\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n}^{(b)},Z_{1:n}^{\prime(b)})\right|-\xi_{n,\ell}\geq\sqrt{\widehat{T}}-\xi_{n}^{\lor}\;\right|\;\mathcal{A}_{1:n},\sqrt{\widehat{T}}>\xi_{n}^{\lor}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\frac{\left(\sqrt{\widehat{T}}-\xi_{n}^{\lor}\right)^{2}}{3a_{\ell}+\sqrt{\widehat{T}}-\xi_{n}^{\lor}}\right)\right\}$
$\leq\min\left\{1,\sum_{\ell=1}^{L}\exp\left(-\frac{3}{2}\sqrt{\widehat{T}}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a_{\ell}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\sqrt{\widehat{T}}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2}\sqrt{\frac{1}{L}\sum_{\ell=1}^{L}\widehat{\mathrm{MMD}}^{2}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{(0)})}+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)\right\}$
$\leq\min\left\{1,L\exp\left(-\frac{3}{2L}\sum_{\ell=1}^{L}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{(0)})\right|+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)\right\}\;.$
Applying a union bound with (23) and the bounds on $a^{\lor}$ (25) and $\xi_{n}^{\lor}$ (26), for any $\eta>0$, with probability at least $\max\{0,1-Le^{-\eta}\}$ over $Z_{1:n}$ and $Z_{1:n}^{(0)}$,
$\exp\left(-\frac{3}{2L}\sum_{\ell=1}^{L}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{(0)})\right|+\frac{3}{2}\xi_{n}^{\lor}+\frac{9}{2}a^{\lor}\right)$
$\quad\leq\exp\left\{-\frac{3}{2L}\sum_{\ell=1}^{L}\left(\Delta_{n,\ell}-2\kappa_{\ell}\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right)\right$.
$\quad\qquad+\left.\frac{3}{2}\left(2\kappa_{\lor}\sqrt{\frac{2}{n}}\right)+\frac{9}{2}\kappa_{\lor}^{2}\left(\frac{2}{n}+16\sqrt{\frac{2}{n^{3}}}\right)\right\}$
$\quad\leq\exp\left\{-\frac{3}{2}\bar{\Delta}_{n}+\kappa_{\lor}\left(6\sqrt{\frac{2}{n}}+3\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{n}\right)\right$.
$\quad\qquad\left.+\kappa_{\lor}^{2}\left(\frac{9}{n}+72\sqrt{\frac{2}{n^{3}}}\right)\right\}$
$\quad=\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+C_{n}^{\lor}(\eta)\right)\;.$
Hence,
$p^{*}\leq\min\left\{1,L\exp\left(-\frac{3}{2}\bar{\Delta}_{n}+C_{n}^{\lor}(\eta)\right)\right\}=:p_{\eta}(\bar{\Delta}_{n})\;.$
For the indicator term in the upper bound of p(T^)p($\widehat${T}), by (23) and (26), with the same probability,
$\mathbbm{1}\left\{\sqrt{\widehat{T}}\leq\xi_{n}^{\lor}\right\}$ $\leq\mathbbm{1}\left\{\sqrt{\frac{1}{L}\sum_{\ell=1}^{L}\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:n},Z_{1:n}^{(0)})}\leq\xi_{n}^{\lor}\right\}$
$\leq\mathbbm{1}\left\{\frac{1}{L}\sum_{\ell=1}^{L}\left|\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}(Z_{1:n},Z_{1:n}^{(0)})\right|\leq\xi_{n}^{\lor}\right\}$
$\leq\mathbbm{1}\left\{\frac{1}{L}\sum_{\ell=1}^{L}\left(\Delta_{n,\ell}-2\kappa_{\ell}\left(\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right)\leq\xi_{n}^{\lor}\right\}$
$\leq\mathbbm{1}\left\{\bar{\Delta}_{n}\leq 2\kappa_{\lor}\left(2\sqrt{\frac{2}{n}}+\sqrt{\eta}\sqrt{\frac{1}{n}+8\sqrt{\frac{2}{n^{3}}}}+\frac{4\eta}{3n}\right)\right\}\;,$
which is zero under assumption (24). Thus, $p(\widehat{T})\leq p_{\eta}(\bar{\Delta}_{n})$ with probability at least $\max\{0,1-Le^{-\eta}\}$.
The rest of the proof is analogous to the second part of Theorem 5.1.
@@
$\Delta_{n}^{\lor}\coloneqq\sup_{\ell\in[L]}\Delta_{n,\ell}=\sup_{\ell\in[L]}\sqrt{\mathbb{E}\left[\left.\widehat{\mathrm{MMD}}_{v,\sigma_{\ell}}^{2}(Z_{1:n},Z_{1:n}^{(0)})\;\right|\;S_{1:n}\right]}$